OpenStudy (anonymous):

Let f(x)= -5x-4 and g(x)=6x-7. find f(x)+g(x)

OpenStudy (goformit100):

-5x-4 + 6x-7 = ?

OpenStudy (luigi0210):

Combine the terms, so: \[\LARGE [f(x)+g(x)]=(-5x-4)+(6x-7)\] \[\LARGE =(-5x+6x)+(-4-7)\] Finish it~

OpenStudy (the_fizicx99):

Like a star!!! ^

OpenStudy (anonymous):

x-11 ?

OpenStudy (luigi0210):

Yup!